OpenStudy (anonymous):

The sum of the interior angles of a heptagon is? A.) 900° B.) 1,080° C.) 1,260° D.) 1,440° I got A, but not for sure.

OpenStudy (mathstudent55):

How many sides does a heptagon have?

OpenStudy (anonymous):

7?

OpenStudy (mathstudent55):

triangle = 3 sides quadrilateral = 4 sides pentagon = 5 sides hexagon = 6 sides heptagon = 7 sides octagon = 9 sides nonagon = 9 sides decagon = 10 sides

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

Now just use the same formula you have used in the last two problems. Here it is again: \(S = (n - 2)180\) where S = sum of the measures of the interior angles, and n = number of sides.

OpenStudy (mathstudent55):

Rewrite the formula exactly as it is. The more you write it, the faster you'll memorize it. Then rewrite it with 7 instead of n. \(S = (n - 2)180\) \(S = (7 - 2)180\) Now do 7 - 2. What is 7 - 2 = ?

OpenStudy (anonymous):

5

OpenStudy (mathstudent55):

Great. Now you write: \(S = 5(180) \) Finally, what is 5 * 180 = ?

OpenStudy (anonymous):

900

OpenStudy (anonymous):

Thank you!

OpenStudy (mathstudent55):

Correct. This is what your notebook should look like: \(S = (n - 2)180\) \(S = (7 - 2)180\) \(S = 5(180) \) \(S = 900\) The sum of the measures of the interior angles of a heptagon is 900 degrees.

OpenStudy (mathstudent55):

If you always start with the formula you will have to advantages: 1. you will get the formula right 2. you will memorize the formula from writing it many times

OpenStudy (mathstudent55):

You're welcome.